Given the linear equation:
- Find the y-coordinates of the given points. Show your work.
- Plot those points and graph the linear equation. Submit your graph through the dropbox.
y = -2x + 5
(0, ), (1, ), (2, )

Answer by richwmiller(9143)

y = -2x + 5
plug in 0 for x and find y
y = -2*(0) + 5
y = -2*(1) + 5
y = -2*(2) + 5